### v3

# A New Upper Bound for Distributed Hypothesis Testing Using the Auxiliary Receiver Approach

Abstract
This paper employs the add-and-subtract technique of the auxiliary receiver approach to establish a new upper bound for the distributed hypothesis testing problem. This new bound has fewer assumptions than the upper bound proposed by Rahman and Wagner, is at least as tight as the bound by Rahman and Wagner, and can outperform it in certain Gaussian settings. Conceptually speaking, unlike Rahman and Wagner, who view their additional receiver as side information, we view it as an auxiliary receiver and use a different manipulation for single-letterization. [^1] This work was supported by the CUHK Direct Grant 4055193.
I Introduction
This paper concerns the problem of hypothesis testing in a distributed setting. Different variants of the problem are defined in the literature (see [6, 1, 2, 3, 4, 5]). In this paper, we consider the simple two-terminal setup as described in Fig. 1. Let $X$ and $Y$ be two sources with discrete alphabets $\mathcal{X}$ and $\mathcal{Y}$, respectively. Two terminals observe i.i.d. sequences $X^{n}$ and $Y^{n}$ over $n$ time instances. Under the null hypothesis $\mathcal{H}_{0}$, the joint distribution is $P_{XY}$, while under the alternative hypothesis $\mathcal{H}_{1}$, it is $Q_{XY}$. The observed sequence of $X$ is transmitted over an $(n,R)$ channel using an encoding function $f_{n}:\mathcal{X}^{n}\rightarrow\mathcal{M}=\{1,2,\ldots,2^{\lfloor nR\rfloor}\}$. The detector receives the message $m=f_{n}(x^{n})$ and decides between the hypotheses $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ based on the following discriminant rule:
$g_{n}(m,y^{n})=\begin{cases}\mathcal{H}_{0},&\text{if }(m,y^{n})\in\mathcal{A}\\ \mathcal{H}_{1},&\text{otherwise}\end{cases}$
where $\mathcal{A}\subseteq\mathcal{M}\times\mathcal{Y}^{n}$ is the acceptance region for $\mathcal{H}_{0}$. See Fig. 1 for an illustration.
The performance of a Distributed Hypothesis Test (DHT) can be quantified by the error exponent with respect to the communication rate $R$. Let $p(\epsilon,n,R)$ denote the smallest achievable error probability under the alternative hypothesis $\mathcal{H}_{1}$ (Type II error) when the source blocklength is $n$, the message length does not exceed $\lfloor nR\rfloor$ bits, and the error probability under the null hypothesis $\mathcal{H}_{0}$ (Type I error) is upper bounded by $\epsilon$. The error exponent of $p(\epsilon,n,R)$ is defined as
$E_{P_{XY},Q_{XY}}(R)\overset{\triangle}{=}\lim_{\epsilon\rightarrow 0}\lim_{n\rightarrow\infty}-\frac{1}{n}\log p(\epsilon,n,R).$
I-A Literature Review
[…]
Figure 1: The problem setup
The following multi-letter characterization of $E_{P_{XY},Q_{XY}}(R)$ is known:
Theorem 1 (Theorems 1 and 4 in [7]).
We have
$E_{P_{XY},Q_{XY}}(R)=\sup_{n,P_{M|X^{n}}}\frac{1}{n}D(P_{MY^{n}}\|Q_{MY^{n}}),$
where the message $M$ takes values in =$\mathcal{M}=\{1,2,\cdots,2^{\lfloor nR\rfloor}\}$.
Quantization-based lower bounds were obtained by Ahlswede and Csiszar [7]:
Theorem 2 (Theorem 5 in [7]).
For any $R>0$, we have
$E_{P_{XY},Q_{XY}}(R)$
$\geq\max_{P_{U|X}:\ I_{P}(U;X)\leq R}\left\{D(P_{X}\|Q_{X})+D(P_{UY}\|Q_{UY})\right\}.$
The use of $P_{U|X}$ implies a Markov chain $U\rightarrow X\rightarrow Y$. The lower bound is not generally tight, as mentioned in [7, Section 3]. The work is then extended and improved by Han [8], and Shimokawa, Han, and Amari (SHA) using the quantization-and-binning technique [9]. Further improvements following this line include [10], [11], [12], [13], and [14].
In terms of the upper bound, the following facts and results are known: The upper bound given by the Neyman-Pearson Lemma in non-distributed hypothesis testing provides the following trivial but useful benchmark at high rates. This bound is called the centralized bound:
$E_{P_{XY},Q_{XY}}(R)\leq D(P_{XY}\|Q_{XY}).$
Conversely, at zero rates, Shalaby and Papamacrou [15] demonstrated that transmitting the sequence type is essentially optimal. The authors in [7] established optimality for the special case of testing against independence where $Q_{XY}=P_{X}P_{Y}$. Rahman and Wagner [16] expanded on this by finding the optimal exponent for “testing against conditional independence" setting. Relevant work also includes [17], and [18]. Authors in [16] obtained the following upper bound:
Definition 1 (Eqns (C14) and (C15) in [16]).
Let $\mathcal{R}(P_{XY},Q_{XY})$ be the set of all $(P_{Z|XY},Q_{Z|XY})$ such that
$Q_{YZ|X}$ $=Q_{Z|X}Q_{Y|Z},$ (1)
$P_{XZ}$ $=Q_{XZ}.$ (2)
Let $\mathcal{S}(P_{XYZ})$ to be the set of all $P_{U|X}$ such that $P_{UXYZ}=P_{U|X}P_{XYZ}$ satisfies
$I_{P}(X;U|Z)\leq R.$
Theorem 3 (Theorem 3, Corollary 5 in [16]).
We have
$E_{P_{XY},Q_{XY}}(R)\leq$ $\min_{P_{Z|XY},Q_{Z|XY}\in\mathcal{R}(P_{XY},Q_{XY})}\{D(P_{YZ}\|Q_{YZ})$
$\qquad+\max_{P_{U|X}\in\mathcal{S}(P_{XYZ})}I_{P}(Y;U|Z)\}.$
Rahman and Wagner interpret $Z$ as side information by employing the following inequality:
$D(P_{MY^{n}}\|Q_{MY^{n}})\leq D(P_{MY^{n}Z^{n}}\|Q_{MY^{n}Z^{n}})$
where $M$ is the transmitted message. In our derivation, we bypass this step and instead utilize a telescopic sum expansion. Consequently, $Z$ does not function as side information but rather as a general auxiliary receiver.
Another upper bound by Hadar et al. [19] involves testing for correlation between Gaussian sequences using the Ornstein-Uhlenbeck process and optimal transport techniques. Notably, neither of the bounds in [16] and [19] is superior to the other one across all parameters for the Gaussian problem.
A more recent improvement is obtained in [20] by giving the detector not just one but a series of side information ( in the sense of [16]). This idea is used in [20] to improve over the bound in [16]. The idea of using multiple auxiliary receivers (instead of just one auxiliary receiver) is independent of the idea discussed in this paper; indeed, our approach can be combined with the idea in [20] to potentially develop even better bounds.
In this paper, we propose a computable upper bound on $E_{P_{XY},Q_{XY}}(R)$ using a systematic manipulation on the auxiliary receivers [21]. The auxiliary receiver approach is a technique for writing new outer bounds for different communication problems. It has been used in [21] to derive new bounds for the broadcast, relay, and interference channels. In this paper, we utilize the “add-and-subtract" technique of the auxiliary receiver approach to derive a new upper bound for the distributed hypothesis testing problem. [^2] The message augmentation technique of the auxiliary receiver approach appears to yield only the bound presented in Proposition 1, which can be proven more readily through a direct method.
I-B Notation
We adopt the following standard notation: for any sequence $(X_{1},X_{2},\cdots,X_{n})$ of random variables, we define
$X^{n}=(X_{1},X_{2},\ldots,X_{n}),$
$X^{i}=(X_{1},\ldots,X_{i}),$
$X_{i+1}^{n}=(X_{i+1},X_{i+2},\ldots,X_{n}),$
$X_{\sim i}=(X_{1},\ldots,X_{i-1},X_{i+1},\ldots,X_{n}).$
For two distributions $\hat{Q}_{X}$ and $Q_{X}$, we say that $\hat{Q}_{X}\ll Q_{X}$ if $\hat{Q}_{X}(x)=0$ whenever $Q_{X}(x)=0$.
All the logs in this paper are in base $e$. Similarly, all the entropies are measured in nats.
II Upper bound using the add-and-subtract technique
II-A Review of the add-and-subtract technique
The add-and-subtract technique works as follows:
  1. 1.
    Begin with a multi-letter expression that can be utilized to derive single-letter upper bounds on capacity. For example, in the point-to-point communication setting, when transmitting a message $M$ over a channel $p(y|x)$, we can consider $I(M;Y^{n})$ as a multi-letter proxy for capacity. In the case of a relay channel $p(y,y_{r}|x,x_{r})$, where $x_{r}$ and $y_{r}$ represent the relay’s input and output, we can treat $I(M;Y^{n})$ or $I(M;Y^{n}Y_{r}^{n})$ as multi-letter expressions serving as upper bounds on the communication rate.
  2. 2.
    Introduce a new auxiliary receiver, denoted as $Z$. Consider a new expression by substituting the auxiliary receiver $Z$ for one of the original receivers. For instance, we can consider $I(M;Z^{n})$ or $I(M;Z^{n}Y_{r}^{n})$ instead of $I(M;Y^{n})$ or $I(M;Y^{n}Y_{r}^{n})$. We call this term the substituted term.
  3. 3.
    Add and subtract the substituted term from the original multi-letter expression. For instance, if $I(M;Y^{n})$ is the original expression and $I(M;Z^{n})$ is the substituted term, we can write:
    $\underbrace{I(M;Y^{n})}_{\text{original term}}=\underbrace{I(M;Y^{n})-I(M;Z^{n})}_{\text{difference term}}+\underbrace{I(M;Z^{n})}_{\text{substituted term}}.$
  4. 4.
    Single-letterize the “difference term" using a telescopic sum expansion. For instance, $I(M;Y^{n})-I(M;Z^{n})$ may be single-letterized as follows:
    $I(M;Y^{n})-I(M;Z^{n})$
    $=\sum_{i}\left(I(M;Y^{i}Z_{i+1}^{n})-I(M;Y^{i-1}Z_{i}^{n})\right).$
  5. 5.
    Lastly, reduce the “substituted term" to a single-letter form by following the procedure of an existing upper bound. This is done by mimicking the bound while formally replacing a legitimate receiver with the auxiliary receiver. For example, in the context of a relay channel, consider the cut-set bound and substitute the original receiver $Y$ with the auxiliary receiver $Z$ throughout all the steps of the bound.
This method produces novel upper bounds for relay, broadcast, and interference channels.
A formal proof of applying the add-and-subtract technique to the distributed hypothesis testing problem is presented in Section II-B. In the remainder of this section, we provide an intuitive discussion of the concept.
For the distributed hypothesis testing problem, we consider $D(P_{MY^{n}}\|Q_{MY^{n}})$ as our multi-letter expression (see Theorem 1). After replacing $Y$ with an auxiliary receiver $Z$, we obtain the expression $D(P_{MZ^{n}}\|Q_{MZ^{n}})$ as our “substituted term". Finally, we form and expand the difference term $D(P_{MY^{n}}\|Q_{MY^{n}})-D(P_{MZ^{n}}\|Q_{MZ^{n}})$ via a telescopic sum as follows:
$\frac{1}{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}D(P_{MY^{n}}\|Q_{MY^{n}})-\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}$
$=\frac{1}{n}\sum_{i=1}^{n}\{D(P_{MY_{1:i-1}Y_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Y_{i}Z_{i+1:n}})$ (3)
$-D(P_{MY_{1:i-1}Z_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Z_{i}Z_{i+1:n}})\}$
$=\frac{1}{n}\sum_{i=1}^{n}\left\{D(P_{U_{i}Y_{i}}\|Q_{U_{i}Y_{i}})-D(P_{U_{i}Z_{i}}\|Q_{U_{i}Z_{i}})\right\}$
where $U_{i}=MY_{1:i-1}Z_{i+1:n}$. With this identification for $U_{i}$, one can also show that $\sum I(U_{i};X_{i}|Y_{i})\leq nR$ and $\sum I(U_{i};X_{i}|Z_{i})\leq nR$. Using the above equations, we can show that the difference term is less than or equal to
$\sup_{P_{U|X},Q_{U|X}}D(P_{UY}\|Q_{UY})-D(P_{UZ}\|Q_{UZ}),$
where $P_{U|X}$ and $Q_{U|X}$ satisfy $I_{P}(U;X|Y)\leq R$, $I_{P}(U;X|Z)\leq R$, $I_{Q}(U;X|Y)\leq R$ and $I_{Q}(U;X|Z)\leq R$. We further simplify this expression by dropping the constraints $I_{Q}(U;X|Y)\leq R$ and $I_{Q}(U;X|Z)\leq R$, and by bounding the objective function from above. This allows us to obtain a simpler expression for comparison with previously known bounds. In Section II-B, we define $G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)$ to as an upper bound on the difference term. This completes Step 4 of the above procedure and yields:
$\underbrace{\frac{1}{n}D(P_{MY^{n}}\|Q_{MY^{n}})}_{\text{originalterm}}$
$=\underbrace{\frac{1}{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}D(P_{MY^{n}}\|Q_{MY^{n}})-\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}}_{\text{difference term}}$
$+\underbrace{\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}_{\text{substituted term}}$
$\leq G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)+\underbrace{\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}_{\text{substituted term}}$
We do not perform Step 5 in this paper and will state the bound simply as follows:
$E_{P_{XY},Q_{XY}}(R)\leq G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)+E_{P_{XZ},Q_{XZ}}(R).$
Indeed, any upper bound on $E_{P_{XZ},Q_{XZ}}(R)$ can be used to derive an upper bound on $E_{P_{XY},Q_{XY}}(R)$. In particular, using the centralized bound $E_{P_{XZ},Q_{XZ}}(R)\leq D(P_{XZ}\|Q_{XZ})$, we obtain
$E_{P_{XY},Q_{XY}}(R)\leq G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)+D(P_{XZ}\|Q_{XZ}).$
However, it is not clear if using the centralized upper bound is the best choice.
Remark 1.
It is useful to compare the distributed hypothesis testing problem with the source model problem in the key agreement context (where the choice of upper bound in Step 5 is clear). We refer the reader to Chapter 22 of [25] for the definition of the key rate capacity $S(X;Y\|Z)$. Given a source $P_{XYZ}$ and any auxiliary receiver $J$, we can show that:
$S(X;Y\|Z)\leq\Delta(P_{XYZ},P_{XYJ})+S(X;Y\|J),$
where
$\Delta(P_{XYZ},P_{XYJ})=\max_{P_{UV|XY}}I(U;Z|V)-I(U;J|V).$
The term $\Delta(P_{XYZ},P_{XYJ})$ serves as an upper bound on the “difference term", and can be obtained through a telescopic sum expansion. In the case of the key agreement problem, we can complete Step 5. Notably, if we adopt the simple upper bound $S(X;Y\|J)\leq I(X;Y|J)$, we arrive at the following upper bound for $S(X;Y|Z)$, which is currently the best known upper bound for this problem [24]:
$S(X;Y\|Z)\leq\Delta(P_{XYZ},P_{XYJ})+I(X;Y|J).$
II-B The bound obtained using the add-and-subtract technique
Definition 2.
Let $\mathcal{T}(P_{XY},P_{XZ})$ to be the set of all $P_{U|X}$ such that
$R\geq\max(I_{P}(U;X|Z),I_{P}(U;X|Y))$
where the mutual information terms $I_{P}(U;X|Z),I_{P}(U;X|Y)$ are computed according to $P_{U|X}P_{XZ}$ and $P_{U|X}P_{XY}$ respectively.
Definition 3.
For any $P_{X}$, $Q_{X}$, $P_{Y|X}$, $P_{Z|X}$, $Q_{Y|X}$ and $Q_{Z|X}$, we define
$f_{P_{Y|X},P_{Z|X},Q_{Y|X},Q_{Z|X}}(P_{X})=\max_{\hat{Q}_{X}}D(P_{Y}\|\hat{Q}_{Y})-D(P_{Z}\|\hat{Q}_{Z})$
where the maximum is over all $\hat{Q}_{X}\ll Q_{X}$ and
$\hat{Q}_{Y}(y)=\sum_{x}\hat{Q}_{X}(x)Q_{Y|X}(y|x),$
$\hat{Q}_{Z}(z)=\sum_{x}\hat{Q}_{X}(x)Q_{Z|X}(z|x).$
Next, let
$G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)$
$=\max_{P_{U|X}\in\mathcal{T}(P_{XY},P_{XZ})}\sum_{u}P(u)f(P_{X|U}(\cdot|u)).$ (4)
Remark 2.
Using the standard application of the Caratheodory theorem, in equation (4), it suffices to restrict to $|\mathcal{U}|\leq|\mathcal{X}|+2$.
Theorem 4.
Assume that $P_{XYZ}$ and $Q_{XYZ}$, satisfy $D(P_{XY}\|Q_{XY})<\infty$ and $D(P_{XZ}\|Q_{XZ})<\infty$. Then, we have
$E_{P_{XY},Q_{XY}}(R)\leq G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)+E_{P_{XZ},Q_{XZ}}(R)$
Proof.
For a given $n$, we assume that $(X^{n},Y^{n},Z^{n})$ are i.i.d. under the two hypothesis:
$P_{X^{n},Y^{n},Z^{n}}=\prod_{i=1}^{n}P_{X_{i},Y_{i},Z_{i}},$
$Q_{X^{n},Y^{n},Z^{n}}=\prod_{i=1}^{n}Q_{X_{i},Y_{i},Z_{i}}.$
Given some encoding map from $X^{n}$ to $M$, we assume
$P_{M,X^{n},Y^{n},Z^{n}}=P_{M|X^{n}}P_{X^{n},Y^{n},Z^{n}},$
$Q_{M,X^{n},Y^{n},Z^{n}}=Q_{M|X^{n}}Q_{X^{n},Y^{n},Z^{n}}.$
We have $P_{M|X^{n}}=Q_{M|X^{n}}$ as the encoding map $f_{n}(\cdot)$ is the same under the two hypotheses.
Take some $\epsilon>0$. Let $n$ and $P_{M|X^{n}}$ be such that
$E_{P_{XY},Q_{XY}}(R)\leq\frac{1}{n}D(P_{MY^{n}}\|Q_{MY^{n}})+\epsilon.$
Next, we utilize the add-and-subtract technique:
$\frac{1}{n}D(P_{MY^{n}}\|Q_{MY^{n}})$
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{n}D(P_{MY^{n}}\|Q_{MY^{n}})-\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}$
$+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}$
From Theorem 1, we have
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{n}D(P_{MZ^{n}}\|Q_{MZ^{n}})}\leq E_{P_{XZ},Q_{XZ}}(R).$
Next, we use the telescopic-sum idea of the add-and-subtract technique to expand the difference term:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}D(P_{MY^{n}}\|Q_{MY^{n}})-D(P_{MZ^{n}}\|Q_{MZ^{n}})}$ (5)
$=\sum_{i=1}^{n}\{D(P_{MY_{1:i-1}Y_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Y_{i}Z_{i+1:n}})$
$-D(P_{MY_{1:i-1}Z_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Z_{i}Z_{i+1:n}})\}$ (6)
where $D(P_{MY^{i}Z_{i+1}^{n}}\|Q_{MY^{i}Z_{i+1}^{n}})$ is finite for every $i$ because
$D(P_{MY^{i}Z_{i+1}^{n}}\|Q_{MY^{i}Z_{i+1}^{n}})$
$\leq D(P_{X^{n}Y^{i}Z_{i+1}^{n}}\|Q_{X^{n}Y^{i}Z_{i+1}^{n}})$
$=iD(P_{XY}\|Q_{XY})+(n-i)D(P_{XZ}\|Q_{XZ})$
<<$\infty$.< ∞ .
Let
$U_{i}=(M,Y_{1:i-1},Z_{i+1:n}).$
Let $T\in\{1,2,\cdots,n\}$ be a time-sharing random variable, independent of all previously defined variables and set
$U=(T,U_{T}),\quad X=X_{T},\quad Y=Y_{T},\quad Z=Z_{T}.$
Observe that
$\sum_{i=1}^{n}\{D(P_{MY_{1:i-1}Y_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Y_{i}Z_{i+1:n}})$
$-D(P_{MY_{1:i-1}Z_{i}Z_{i+1:n}}\|Q_{MY_{1:i-1}Z_{i}Z_{i+1:n}})\}$
$=\sum_{i=1}^{n}D(P_{U_{i}Y_{i}}\|Q_{U_{i}Y_{i}})-D(P_{U_{i}Z_{i}}\|Q_{U_{i}Z_{i}})$
$=\sum_{i=1}^{n}\underbrace{D(P_{U_{i}}\|Q_{U_{i}})}_{\text{cancelled}}+D(P_{Y_{i}|U_{i}}\|Q_{Y_{i}|U_{i}}|P_{U_{i}})$
$-\underbrace{D(P_{U_{i}}\|Q_{U_{i}})}_{\text{cancelled}}-D(P_{Z_{i}|U_{i}}\|Q_{Z_{i}|U_{i}}|P_{U_{i}}).$
Observe that under either $P_{M,X^{n},Y^{n},Z^{n}}$ or $Q_{M,X^{n},Y^{n},Z^{n}}$, we have the following Markov chain:
$U_{i}\rightarrow X_{i}\rightarrow Y_{i}Z_{i}$
as $00\leq I(MY_{1:i-1}Z_{i+1:n};Y_{i}Z_{i}|X_{i})\leq I(X^{n}Y_{1:i-1}Z_{i+1:n};Y_{i}Z_{i}|X_{i})=$ 0 by the i.i.d assumption. Then
$\sum_{i=1}^{n}D(P_{Y_{i}|U_{i}}\|Q_{Y_{i}\|U_{i}}|P_{U_{i}})-D(P_{Z_{i}|U_{i}}\|Q_{Z_{i}|U_{i}}|P_{U_{i}})$
$=\sum_{i=1}^{n}\sum_{u_{i}}P_{U_{i}}(u_{i})\left[D(P_{Y_{i}|u_{i}}\|Q_{Y_{i}|u_{i}})-D(P_{Z_{i}|u_{i}}\|Q_{Z_{i}|u_{i}})\right]$
$=n\cdot\sum_{u}P_{U}(u)\left[D(P_{Y|u}\|Q_{Y|u})-D(P_{Z|u}\|Q_{Z|u})\right]$
$\leq n\cdot\sum_{u}P_{U}(u)\max_{\hat{Q}_{X}}\left[D(P_{Y|u}\|\hat{Q}_{Y})-D(P_{Z|u}\|\hat{Q}_{Z})\right]$
where the maximum is over $\hat{Q}_{X}\ll Q_{X}$ (observe that if $Q_{X}(x)=0$, then $Q_{X|u}(x)=0$).
Next, we have $I_{P}(U;X|Z)\leq R$ since
$nR$ $\geq H_{P}(M)$ (7)
$\geq I_{P}(M;X^{n}|Z^{n})$ (8)
$=\sum_{i=1}^{n}I_{P}(M;X_{i}|Z^{n}X^{i-1})$
$=\sum_{i=1}^{n}I_{P}(MX^{i-1}Z_{\sim i};X_{i}|Z_{i})$ (9)
$=\sum_{i=1}^{n}I_{P}(MX^{i-1}Y^{i-1}Z_{\sim i};X_{i}|Z_{i})$ (10)
$\geq\sum_{i=1}^{n}I_{P}(MY^{i-1}Z_{i+1}^{n};X_{i}|Z_{i})$
$=\sum_{i=1}^{n}I_{P}(U_{i};X_{i}|Z_{i})$
$=n\cdot I_{P}(U;X|Z)$
where (9) follows from the i.i.d assumption, and (10) follows from $I(Y^{i-1};X_{i}|Z^{n}X^{i-1}M)=0$ as
$I(Y^{i-1};X_{i}|Z^{n}X^{i-1}M)$ $\leq I(Y^{i-1};X_{i}M|Z^{n}X^{i-1})$
$\leq I(Y^{i-1};X^{n}|Z^{n}X^{i-1})$
$=I(Y^{i-1};X^{i}|Z^{n})=0.$
The proof for $R\geq I_{P}(U;X|Y)$ is similar. This completes the proof. ∎
The following theorem yields an explicit upper bound in a special case, which is useful in comparing our bound with the previously known bounds:
Theorem 5.
Take an arbitrary pair $(P_{Z|XY},Q_{Z|XY})$ such that $Q_{YZ|X}=Q_{Z|X}Q_{Y|Z}$. Then for any $P_{UX}$ we have
$\sum_{u}P(u)f(P_{X|U}(\cdot|u))\leq D(P_{YZU}\|P_{UZ}Q_{Y|Z}).$ (11)
Proof.
Given some $\hat{Q}_{X}$, let
$\hat{Q}_{XYZ}=\hat{Q}_{X}Q_{YZ|X}=\hat{Q}_{X}Q_{Z|X}Q_{Y|Z}.$
Then, we have
$\sum_{u}P(u)\max_{\hat{Q}_{X}}D(P_{Y|u}\|\hat{Q}_{Y})-D(P_{Z|u}\|\hat{Q}_{Z})$
$\leq\sum_{u}P(u)\max_{\hat{Q}_{X}}D(P_{YZ|u}\|\hat{Q}_{YZ})-D(P_{Z|u}\|\hat{Q}_{Z})$ (12)
$=\sum_{u}P(u)\max_{\hat{Q}_{X}}\sum_{z}P(z|u)D(P_{Y|z,u}\|\hat{Q}_{Y|z})$
$=\sum_{u}P(u)\max_{\hat{Q}_{X}}\sum_{z}P(z|u)D(P_{Y|z,u}\|Q_{Y|z})$ (13)
$=\sum_{u}P(u)\sum_{z}P(z|u)D(P_{Y|z,u}\|Q_{Y|z})$
$=D(P_{YZU}\|P_{UZ}Q_{Y|Z})$
where (12) follows from by the data processing inequality, and (13) follows from the definition of $\hat{Q}_{XYZ}$. ∎
We remark that, unlike the left-hand side, the right-hand side in (11) does not depend on $Q_{XZ}$, so one may assume $Q_{XZ}=P_{XZ}$ if using Theorem 5.
III Comparison with previous bounds
Next, we show that the bound in Theorem 4 improves over the bounds given in [20] and [16].
Lemma 1.
Given any $(P_{Z|XY},Q_{Z|XY})\in\mathcal{R}(P_{XY},Q_{XY})$ as defined in Definition 1 and for every $P_{X|U}$, we have
$\sum_{u}P(u)f(P_{X|U}(\cdot|u))\leq I_{P}(Y;U|Z)+D(P_{YZ}\|Q_{YZ}).$ (14)
Consequently,
$G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)$
$\leq\max_{P_{U|X}\in\mathcal{T}(P_{XY},P_{XZ})}I_{P}(Y;U|Z)+D(P_{YZ}\|Q_{YZ}).$
Proof.
From the definition of $\mathcal{R}(P_{XY},Q_{XY})$ we obtain
$P_{Z}=Q_{Z}$
Then immediately,
$D(P_{YZU}\|P_{UZ}Q_{Y|Z})$
$=\sum_{u,y,z}P_{YZU}\log\Big{(}\frac{P_{YZU}Q_{Z}}{P_{UZ}Q_{YZ}}\Big{)}$
$=\sum_{u,y,z}P_{YZU}\log\Big{(}\frac{P_{YU|Z}}{P_{U|Z}P_{Y|Z}}\cdot\frac{P_{YZ}}{Q_{YZ}}\Big{)}$
$=I_{P}(Y;U|Z)+D(P_{YZ}\|Q_{YZ}).$
Apply Theorem 5 completes the proof.
Corollary 1.
Given that $(P_{Z|XY},Q_{Z|XY})\in\mathcal{R}(P_{XY},Q_{XY})$, we have
$E_{P_{XY},Q_{XY}}(R)\leq\max_{P_{U|X}\in\mathcal{T}(P_{XY},P_{XZ})}I_{P}(Y;U|Z)+D(P_{YZ}\|Q_{YZ}).$
This corollary follows from $E_{P_{XZ},Q_{XZ}}(R)=0$ by the assumption $P_{XZ}=Q_{XZ}$.
Note that the bound in Corollary 1 is a weaker version of the auxiliary receiver bound. The bound in Corollary 1 is less than or equal to the bound in [16] since $\mathcal{T}(P_{XY},P_{XZ})\subset\mathcal{S}_{P_{XYZ}}$ as $\mathcal{S}_{P_{XYZ}}$ is the set of $P_{U|X}$ satisfying
$R\geq I_{P}(X;U|Z)$
while $P_{U|X}\in\mathcal{T}(P_{XY},P_{XZ})$ satisfies
$R\geq\max(I_{P}(X;U|Y),I_{P}(X;U|Z)).$
Next, as a sanity check, we consider the example of “test against conditional independence" from [16]:
Example 1 (Test against conditional independence).
Assume that $Y=(J,Z,\hat{Y})$ for some random variables $J$, $Z$ and $\hat{Y}$. We denote the distribution of $J,Z,\hat{Y}$ under distribtion $P$ and $Q$ by $P_{J,Z,\hat{Y}}$ and $Q_{J,Z,\hat{Y}}$. Moreover, assume that $P_{XY}=P_{XJZ\hat{Y}}$ and $Q_{XY}=Q_{XJZ\hat{Y}}$ satisfy the following constraints: $I_{Q}(\hat{Y};XJ|Z)=0$, $P_{XJZ}=Q_{XJZ}$ and $P_{\hat{Y}Z}=Q_{\hat{Y}Z}$. Then, the upper bound in Corollary 1 implies the following bound:
$E_{P_{XY},Q_{XY}}(R)\leq\max_{P_{U|X}}I_{P}(\hat{Y};UJ|Z)$
where the maximum is over $P_{U|X}$ satisfying $I_{P}(U;X|ZJ)\leq R$.
Proof of Example 1.
Let $Z^{\prime}=(Z,J)$. Note that (1) and (2) hold for $(X,Y,Z^{\prime})=(X,Y,ZJ)$, i.e.,
$Q_{YZJ|X}$ $=Q_{ZJ|X}Q_{Y|ZJ},$ (15)
$P_{XZJ}$ $=Q_{XZJ},$ (16)
due to our assumptions. Then, $(P_{Z^{\prime}|XY},Q_{Z^{\prime}|XY})\in\mathcal{R}(P_{XY},Q_{XY})$ and we obtain
$E_{P_{XY},Q_{XY}}(R)$
$\leq\max_{P_{U|X}\in\mathcal{T}(P_{XY},P_{XZ^{\prime}})}I_{P}(Y;U|Z^{\prime})+D(P_{YZ^{\prime}}\|Q_{YZ^{\prime}})$
$=\max_{P_{U|X}}I_{P}(Y;U|ZJ)+D(P_{YZJ}\|Q_{YZJ})$
$=\max_{P_{U|X}}I_{P}(Y;U|ZJ)+D(P_{Y}\|Q_{Y})$
where the maximum is over $P_{U|X}$ satisfying $I_{P}(U;X|YJ)\leq R$ and $I_{P}(U;X|ZJ)\leq R$. Observe that $I_{P}(U;X|YJ)\leq I_{P}(U;X|ZJ)$ because $Z$ is a function of $Y$. Thus, $I_{P}(U;X|YJ)\leq R$ is redundant. Further, we have
$I_{P}(Y;U|ZJ)+D(P_{Y}\|Q_{Y})$
$=I_{P}(\hat{Y};U|ZJ)+D(P_{\hat{Y}ZJ}\|Q_{\hat{Y}ZJ})$
$=I_{P}(\hat{Y};U|ZJ)+D(P_{\hat{Y}ZJ}\|P_{\hat{Y}Z}Q_{J|\hat{Y}Z})$
$=I_{P}(\hat{Y};U|ZJ)+D(P_{\hat{Y}ZJ}\|P_{\hat{Y}Z}Q_{J|Z})$
$=I_{P}(\hat{Y};U|ZJ)+D(P_{\hat{Y}ZJ}\|P_{\hat{Y}Z}P_{J|Z})$
$=I_{P}(\hat{Y};U|ZJ)+I_{P}(\hat{Y};J|Z)$
$=I_{P}(\hat{Y};UJ|Z).$
III-A Strict improvement for Gaussian Sources
Next, we show that the bound in Corollary 1 can provide better bounds than the one in [16]. Authors in [16] evaluate their bound for the following Gaussian setting: consider a Gaussian scenario where $X$ and $Y$ are jointly Gaussian sources with zero mean and unit variance. Assume that their correlation coefficients under hypotheses $H_{0}$ and $H_{1}$ are $\rho_{0}$ and $\rho_{1}$, respectively. In other words, under the hypothesis $H_{i}$, we have
$K_{X,Y}\sim\mathcal{N}\bigg{(}0,\begin{bmatrix}1&\rho_{i}\\ \rho_{i}&1\end{bmatrix}\bigg{)}$
where $\rho_{0}\in[-1,1]$, and without loss of generality, $\rho_{1}\in[0,1]$. Even though we stated our bounds for discrete settings, a similar argument goes through for jointly Gaussian random variables. Moreover, in this section, we measure the communication rate in nats per symbol and use the natural logarithm.
III-A1 Existing Bounds
The authors in [16] and [19] provide the following bounds:
Theorem 6 (Theorem 9 in [16]).
For the Gaussian setting, we have
$E_{RW}(R)$ $=\frac{1}{2}\log\left(\frac{1}{1-{\rho}^{2}+{\rho}^{2}e^{-2R}}\right)+C$
where
$\mathcal{D}_{1}$$\triangleq\{(\rho_{0},\rho_{1}):0\leq\rho_{1}<\rho_{0}<1\},$
$\mathcal{D}_{2}$$\triangleq\{(\rho_{0},\rho_{1}):0\leq\rho_{1}\hskip 3.61371pt\textrm{and}\hskip 3.61371pt2\rho_{1}-1\leq\rho_{0}<\rho_{1}\},$
$\mathcal{D}_{3}$$\triangleq\Big{\{}(\rho_{0},\rho_{1}):-1<\rho_{0}\leq 2\rho_{1}-1\hskip 3.61371pt\textrm{and}$
$\hskip 3.61371pt\frac{2\rho_{1}}{1-\rho_{1}}\leq\frac{1}{2}\log\left(\frac{1-\rho_{1}^{2}}{1-\rho_{0}^{2}}\right)-\frac{\rho_{1}(\rho_{0}-\rho_{1})}{1-\rho_{1}^{2}}\Big{\}}.$
$\rho\triangleq\left\{\begin{array}[]{l l}\frac{\rho_{0}-\rho_{1}}{1-\rho_{1}}&\quad\mbox{if $(\rho_{0},\rho_{1})$ is in $\mathcal{D}_{1}\cup\mathcal{D}_{2}$}\\ \frac{\rho_{0}+\rho_{1}}{1-\rho_{1}}&\quad\mbox{if $(\rho_{0},\rho_{1})$ is in$\mathcal{D}_{3}$}.\\ \end{array}\right.$
$C\triangleq\left\{\begin{array}[]{l l}0&\quad\mbox{if $(\rho_{0},\rho_{1})$ is in $\mathcal{D}_{1}\cup\mathcal{D}_{2}$}\\ \frac{2\rho_{1}}{1-\rho_{1}}&\quad\mbox{if $(\rho_{0},\rho_{1})$ is in $\mathcal{D}_{3}$}.\\ \end{array}\right.$
Theorem 7 (Theorem 1 in [19]).
For any $\rho_{0},\rho_{1}\geq 0$,
$E_{HLPS}(R)=\frac{R}{\Big{(}\frac{1-\min\{\rho_{0},\rho_{1}\}}{\rho_{1}-\rho_{0}}\Big{)}^{2}-1}.$
In terms of lower bounds, the expression of SHA exponent in the Gaussian setting was given in [20]:
Theorem 8 (Equations (35)-(40) in [20]).
For the Gaussian hypothesis testing the following exponent is achievable:
$E_{SHA}(R)=\sup_{\eta\geq 0}\max\Big{[}E_{\eta}(R),\max_{0\leq\beta\leq\beta_{\max}}$
$\min\{E_{\eta}(\beta R),\overline{E}_{\eta}(\beta R)-(\beta-1)R+E_{\eta}(\beta R){\mathbf{1}}_{\{\rho_{0}<\rho_{1}\}}\}\Big{]}$
where
$E_{\eta}(R)=D\Big{(}\frac{(\rho_{1}-\eta)^{2}\sigma_{R}^{2}+\rho_{1}^{2}(1-\sigma_{R}^{2})}{1+\eta(\eta-2\rho_{1})\sigma_{R}^{2}},\frac{1+\eta(\eta-2\rho_{0})\sigma_{R}^{2}}{1+\eta(\eta-2\rho_{1})\sigma_{R}^{2}}\Big{)}$
$\overline{E}_{\eta}(R)=D\Big{(}\frac{\eta^{2}\sigma_{R}^{2}}{1+\eta^{2}\sigma_{R}^{2}},\frac{1+\eta(\eta-2\rho_{0})\sigma_{R}^{2}}{1+\eta^{2}\sigma_{R}^{2}}\Big{)}$
$\sigma_{R}^{2}=1-\exp(-2R)$
$D(\gamma,P)=\frac{\gamma+P-A(\gamma,P)}{2(1-\gamma)}-\frac{1}{2}\log\frac{A(\gamma,P)-1+\gamma}{2\gamma}$
$A(\gamma,P)=\sqrt{(1-\gamma)^{2}+4P\gamma}$
The maximization limit $\beta_{\max}$ is the unique solution of $\overline{E}_{\eta}(\beta R)=(\beta-1)R$.
III-A2 New Bound
Below, we state our new bound by optimizing the lowest upper bound using Corollary 1 (which is a weaker version of our main bound in Theorem 4):
Theorem 9.
For the standard Gaussian setting and assuming $\rho_{0}\in[-1,1],\rho_{1}\in[0,1]$, we have
$E_{P_{XY},Q_{XY}}(R)\leq\min_{(P_{XY},Q_{XY})\in\mathcal{R}}$$\Big{\{}\max_{P_{U|X}\in\mathcal{T}_{R}}I_{P}(Y;U|Z)$
$+D(P_{YZ}\|Q_{YZ})\Big{\}},$
where
$\mathcal{R}$ =$=\big{\{}(P_{Z|XY},Q_{Z|XY}):Q_{YZ|X}=Q_{Z|X}Q_{Y|Z},$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad P_{XZ}=Q_{XZ}\big{\}}$
$\mathcal{T}_{R}$ =$=\left\{P_{U|X}:R\geq\max(I_{P}(U;X|Z),I_{P}(U;X|Y)\right\}.$
For evaluation of the bound, one should use the following expression:
$E_{P_{XY},Q_{XY}}(R)\leq\min_{b,g}$$\Big{\{}I_{P}(Y;U|Z)+D(P_{YZ}\|Q_{YZ})\Big{\}},$
where
$\hat{\sigma}^{2}$ $=\max\left(\frac{1-g^{2}}{\exp(2R)-g^{2}},\frac{1-(a+bg)^{2}}{\exp(2R)-(a+bg)^{2}}\right).$
$I_{P}(Y;U|Z)$ $=\frac{1}{2}\log\left\{\frac{1-[(\rho_{0}-bg)g+b]^{2}}{1-\rho_{0}^{2}(1-\hat{\sigma}^{2})-\frac{[\rho_{0}g\hat{\sigma}^{2}+b(1-g^{2})]^{2}}{1-(1-\hat{\sigma}^{2})g^{2}}}\right\},$
$D(P_{YZ}\|Q_{YZ})$ $=\frac{g^{2}-{\rho_{1}}{g}[(\rho_{0}-bg)g+b]}{{g^{2}}-{\rho_{1}^{2}}}-1$
$\qquad\qquad+\frac{1}{2}\log\frac{{g^{2}}-{\rho_{1}^{2}}}{{g^{2}}-{g^{2}}[(\rho_{0}-bg)g+b]^{2}},$
where the minimum is over $b,g$ satisfying
$g\in[-1,1],\quad b^{2}(1-g^{2})\leq 1-\rho_{0}^{2},\quad((\rho_{0}-bg)g+b)^{2}\leq 1,$
$\rho_{1}^{2}\leq g^{2}.$
Plots comparing the bound with all the other known bounds are given in Fig. 2 and Fig. 3. One can observe that our bound is always better than the bound in [16] for these parameters. Computing the bound in [20] is difficult, so we did not replicate the simulation done by [20] and just extracted the points from the curve given in [20]. One can observe that the auxiliary receiver bound is better than the bound in [20] when $\rho_{0}$ is not close to one. For $\rho_{0}$ close to one, the bound in [20] is better, which suggests multiple auxiliary receivers can yield strictly better bounds for large $\rho_{0}$. We have discussed combining the idea in [20] with our technique in Section III-B.
Figure 2: Exponents (normalized by $(\rho_{0}-\rho_{1})^{2}$) as a function of $\rho_{0}$ for fixed $\rho_{1}=0.7$ and rate $R=0.5$. Red is SHA’s achievable exponent. Cyan is the centralized upper bound. Blue is HLPS upper bound in [19]. Magenta is RW upper bound in [16]. Black is our bound in Theorem 9. Dashed black is the upper bound extracted from the plot in [20].
Figure 3: Exponents (normalized by $(\rho_{0}-\rho_{1})^{2}$) with fixed $\rho_{1}=0.2$ and rate $R=0.1$. In this case, the other bounds discussed are large and thus not shown here. $\rho_{0}\in[\rho_{1},1]$ for a better visual illustration with RW upper bound in [16]. AXR (auxiliary receiver) upper bound goes to RW upper bound as $\rho_{0}\rightarrow 1$ and to the achievable SHA exponent as $\rho_{0}\rightarrow\rho_{1}$.
Proof of Theorem 9.
Under $P$, assume that
$Z=gX+\sqrt{1-g^{2}}N_{1},$
$Y=aX+bZ+N_{3}\sqrt{1-a^{2}-b^{2}-2abg},$
and under $Q$, assume that
$Z=gX+\sqrt{1-g^{2}}N_{1},$
$Y=fZ+\sqrt{1-f^{2}}N_{2},$
where $I(N_{1};X)=I(N_{2};Z)=I(N_{3};X,Z)=0$ and $X,Y,Z,N_{1},N_{2},N_{3}$ are standard Gaussian random variables.
Then we remove redundant parameters using $\rho_{0}$ and $\rho_{1}$ and obtain constraints on the free parameters. Under $P$, the covariance of $X$ and $Y$ is $\rho_{0}$:
$Cov(X,Y)=a+bg=\rho_{0},$
$Cov(Y,Z)=ag+b\in[-1,1],$
$Cov(Y,N_{2})=\sqrt{1-(a^{2}+b^{2}+2abg)}\in[-1,1],$
$Cov(X,Z)=g\in[-1,1].$
Under $Q$, the covariance of $X$ and $Y$ is $\rho_{1}$, thus:
$Cov(X,Y)=Cov(X,fgX)=fg=\rho_{1},$
$Cov(Y,Z)=f\in[-1,1].$
The above characterization gives
$I_{P}(Y,U|Z)$
$=\frac{1}{2}\log\left\{\frac{1-(ag+b)^{2}}{1-(a+bg)^{2}(1-\sigma^{2})-\frac{[(a+bg)g\sigma^{2}+b(1-g^{2})]^{2}}{1-(1-\sigma^{2})g^{2}}}\right\},$
and
$D(P_{YZ}\|Q_{YZ})=\frac{1-f(ag+b)}{1-f^{2}}-1+\frac{1}{2}\log\frac{1-f^{2}}{1-(ag+b)^{2}}.$
We show the optimality of a jointly Gaussian distribution on $(U,X)$ when evaluating $G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)$ (and hence the above weaker upper bound) in Appendix A.
If the variance of $P_{X|U}$ is $\sigma^{2}\leq 1$. The derivative of the denominator of $I_{P}(Y,U|Z)$ is $\frac{a^{2}(g^{2}-1)^{2}}{(g^{2}\sigma^{2}-g^{2}+1)^{2}}\geq 0$. The optimal $\sigma^{2}$ that maximizes $I_{P}(Y,U|Z)$ within feasible region $\mathcal{T}_{R}$ is then given by:
$\hat{\sigma}^{2}=\max\left(\frac{1-g^{2}}{\exp(2R)-g^{2}},\frac{1-(a+bg)^{2}}{\exp(2R)-(a+bg)^{2}}\right).$
III-B The bound in [20] and a sequence of auxiliary receivers
Definition 4 ([20]).
[…]
$Q_{YZ|X}$ $=Q_{Z|X}Q_{Y|Z}$ (17)
$P_{Y|Z}$ $=Q_{Y|Z}$ (18)
The following bound was given in [20], which is in the same spirit as the auxiliary receiver bound:
Lemma 2 (Lemma 1 in [20]).
For any $(P_{Z|XY},Q_{Z|XY})$ in ~$\tilde{\mathcal{R}}(P_{XY},Q_{XY})$:
$D(P_{MY^{n}}\|Q_{MY^{n}})$ $\leq\max_{P_{U|X}\in\mathcal{S}(P_{XYZ})}I_{P}(Y;U|Z)$
$+D(P_{MZ^{n}}\|Q_{MZ^{n}})$
Observe that the bound stated in Theorem 4 allows for arbitrary choices of $P_{Z|XY}$ and $Q_{Z|XY}$, while the one in Lemma 2 is restricted to a class of auxiliary receivers $(P_{Z|XY},Q_{Z|XY})$ in ~$\tilde{\mathcal{R}}(P_{XY},Q_{XY})$. Note that $(P_{Z|XY},Q_{Z|XY})$ in ~$\tilde{\mathcal{R}}(P_{XY},Q_{XY})$ requires $P_{Y|Z}=Q_{Y|Z}$. Therefore,
$D(P_{YZU}\|P_{UZ}Q_{Y|Z})$ $=D(P_{YZU}\|P_{UZ}P_{Y|Z})=I_{P}(Y;U|Z).$
Theorem 5 then implies that the bound in Theorem 4 is less than or equal to the one in Lemma 2.
The author in [20] suggests the use of a sequence of auxiliary receivers $Z_{1},Z_{2},\cdots,Z_{k}$ to derive an improved bound. This approach can also be integrated with our method, which might result in further improvement. In other words, for any sequence of auxiliary receivers $Z_{1},Z_{2},\cdots,Z_{k}$ we can write
$E_{P_{XY},Q_{XY}}(R)$ (19)
$\leq G_{P_{XY},Q_{XY},P_{XZ_{1}},Q_{XZ_{1}}}(R)$ (20)
$+\sum_{j=1}^{k-1}G_{P_{XZ_{j}},Q_{XZ_{j}},P_{XZ_{j+1}},Q_{XZ_{j+1}}}(R)+E_{P_{XZ_{k}},Q_{XZ_{k}}}(R).$ (21)
To motivate (21), consider the bound
$E_{P_{XY},Q_{XY}}(R)\leq G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)+E_{P_{XZ},Q_{XZ}}(R).$
Another trivial bound is the centralized bound:
$E_{P_{XY},Q_{XY}}(R)\leq D(P_{XY}\|Q_{XY}).$
The following example shows that the centralized bound might be better in some cases:
Example 2.
Assume that $P_{Y|X},Q_{Y|X},P_{Z|X}$ and $Q_{Z|X}$ are additive Gaussian noise channels with additive noise variances $\sigma_{P,Y|X}^{2}$, $\sigma_{Q,Y|X}^{2}$,$\sigma_{P,Z|X}^{2}$, $\sigma_{Q,Z|X}^{2}$ respectively. Assume that $X\sim\mathcal{N}(0,\sigma_{X}^{2})$ under $P$. We claim that $f_{P_{Y|X},P_{Z|X},Q_{Y|X},Q_{Z|X}}(P_{X})=\infty$ if $\sigma_{Q,Y|X}<\sigma_{Q,Z|X}$. Assume that $X\sim\mathcal{N}(k,\sigma_{X}^{2})$ under $\hat{Q}_{X}$. Let
$\sigma_{P,Y}$ $=\sqrt{\sigma_{X}^{2}+\sigma_{P,Y|X}^{2}},\qquad\sigma_{P,Z}=\sqrt{\sigma_{X}^{2}+\sigma_{P,Z|X}^{2}},$
$\sigma_{Q,Y}$ $=\sqrt{\sigma_{X}^{2}+\sigma_{Q,Y|X}^{2}},\qquad\sigma_{Q,Z}=\sqrt{\sigma_{X}^{2}+\sigma_{Q,Z|X}^{2}}.$
Then, we get
$f_{P_{Y|X},P_{Z|X},Q_{Y|X},Q_{Z|X}}(P_{X})$
$\geq D(P_{Y}\|\hat{Q}_{Y})-D(P_{Z}\|\hat{Q}_{Z})$
$=\log\frac{\sigma_{Q,Y}}{\sigma_{P,Y}}+\frac{\sigma_{P,Y}^{2}-\sigma^{2}_{Q,Y}+k^{2}}{2\sigma^{2}_{Q,Y}}$
$-\log\frac{\sigma_{Q,Z}}{\sigma_{P,Z}}-\frac{\sigma_{P,Z}^{2}-\sigma^{2}_{Q,Z}+k^{2}}{2\sigma^{2}_{Q,Z}}$
Letting $k\rightarrow\infty$, we get that $f_{P_{Y|X},P_{Z|X},Q_{Y|X},Q_{Z|X}}(P_{X})=\infty$.
Proposition 1.
Take some arbitrary $P_{J|XYZ}$ and $Q_{J|XYZ}$. Assume that $P_{XYZJ}$ and $Q_{XYZJ}$, satisfy $D(P_{XYJ}\|Q_{XYJ})<\infty$ and $D(P_{XZJ}\|Q_{XZJ})<\infty$. Let $Y^{\prime}=(Y,J)$ and $Z^{\prime}=(Z,J)$. Then, we have
$E_{P_{XY},Q_{XY}}(R)\leq G_{P_{XY^{\prime}},Q_{XY^{\prime}},P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)$ (22)
$+E_{P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)$ (23)
Proof.
Consider (21) with the choice $Z_{1}=Y^{\prime}$ and $Z_{2}=Z^{\prime}$:
$E_{P_{XY},Q_{XY}}(R)$
$\leq G_{P_{XY},Q_{XY},P_{XY^{\prime}},Q_{XY^{\prime}}}(R)$
$+G_{P_{XY^{\prime}},Q_{XY^{\prime}},P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)+E_{P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)$
$\leq G_{P_{XY^{\prime}},Q_{XY^{\prime}},P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)+E_{P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)$
as $G_{P_{XY},Q_{XY},P_{XY^{\prime}},Q_{XY^{\prime}}}(R)\leq 0$. ∎
Remark 3.
If we set $J$ to be constant, both bounds recover the original bound. If we set $J=X$, we recover the centralized bound because
$f_{P_{Y^{\prime}|X},P_{Z^{\prime}|X},Q_{Y^{\prime}|X},Q_{Z^{\prime}|X}}(P_{X})$
$=\max_{\hat{Q}_{X}}D(P_{XY}\|\hat{Q}_{XY})-D(P_{XZ}\|\hat{Q}_{ZX})$
$=\max_{\hat{Q}_{X}}D(P_{Y|X}\|\hat{Q}_{Y|X}|P_{X})-D(P_{Z|X}\|\hat{Q}_{Z|X}|P_{X})$
$=D(P_{Y|X}\|Q_{Y|X}|P_{X})-D(P_{Z|X}\|Q_{Z|X}|P_{X})$
where we used the fact that $\hat{Q}_{Z|X}=Q_{Z|X}$ and $\hat{Q}_{Y|X}=Q_{Y|X}$. Observe that $f_{P_{Y^{\prime}|X},P_{Z^{\prime}|X},Q_{Y^{\prime}|X},Q_{Z^{\prime}|X}}(P_{X})$ is linear in $P_{X}$, and it is optimal to choose constant $U$ when evaluating $G_{P_{XY^{\prime}},Q_{XY^{\prime}},P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)$. We obtain that
$G_{P_{XY^{\prime}},Q_{XY^{\prime}},P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)$
$=D(P_{Y|X}\|Q_{Y|X}|P_{X})-D(P_{Z|X}\|Q_{Z|X}|P_{X})$
$=D(P_{XY}\|Q_{XY})-D(P_{XZ}\|Q_{XZ}).$
Thus, from (23) and using $E_{P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)=D(P_{XZ}\|Q_{XZ})$ when $J=X$, we obtain
$G_{P_{XY^{\prime}},Q_{XY^{\prime}},P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)+E_{P_{XZ^{\prime}},Q_{XZ^{\prime}}}(R)=D(P_{XY}\|Q_{XY})$
as desired.
Appendix A Evaluation of the upper bound in the Gaussian setting
Consider the following optimization problem
$G_{P_{XY},Q_{XY},P_{XZ},Q_{XZ}}(R)$
$=\sup_{P_{U|X}}\mathbb{E}_{P_{U}}\left[\max_{\hat{Q}_{X}}D(P_{Y|U}\|\hat{Q}_{Y})-D(P_{Z|U}\|\hat{Q}_{Z})\right]$
where the supremum is over $P_{U|X}$ satisfying $I_{P}(U;X|Y)\leq R$ and $I_{P}(U;X|Z)\leq R$. Assume that $X$ is a scalar standard Gaussian random variable under both hypotheses. Assume that $P_{Y|X},Q_{Y|X},P_{Z|X}$ and $Q_{Z|X}$ are additive Gaussian noise channels with additive noise variances $\sigma_{P,Y|X}^{2}$, $\sigma_{Q,Y|X}^{2}$, $\sigma_{P,Z|X}^{2}$, $\sigma_{Q,Z|X}^{2}$ respectively. If $\sigma_{Q,Z|X}>\sigma_{Q,Y|X}$, the above expression becomes infinity. This can be seen by choosing a constant random variable $U$ and following the choice in Example 2 given in Section III-B.
In the following, we assume that $\sigma_{Q,Z|X}\leq\sigma_{Q,Y|X}$. In this case, we can find a jointly Gaussian distribution $Q_{XYZ}$ (a coupling) such that
$Q_{XYZ}=Q_{X}Q_{Z|X}Q_{Y|Z}.$
Note that the assumption $Q_{XYZ}=Q_{X}Q_{Z|X}Q_{Y|Z}$ would then imply that the inner maximum over $\hat{Q}_{X}$ is bounded for every $P_{U|X}$. Since
$\mathbb{E}_{P_{U}}\left[\max_{\hat{Q}_{X}}D(P_{Y|U}\|\hat{Q}_{Y})-D(P_{Z|U}\|\hat{Q}_{Z})\right]$
$=-h_{P}(Y|U)+h_{P}(Z|U)+\mathbb{E}_{P_{U}}\max_{\hat{Q}_{X}}\mathbb{E}_{P_{YZ|U}}\log\frac{\hat{Q}_{Z}}{\hat{Q}_{Y}}$
we can observe that for every $P_{UX}$ we have
$\mathbb{E}_{P_{U}}\max_{\hat{Q}_{X}}\mathbb{E}_{P_{YZ|U}}\log\frac{\hat{Q}_{Z}}{\hat{Q}_{Y}}\leq\mathbb{E}_{P_{U}}\max_{\hat{Q}_{X}}\mathbb{E}_{P_{YZ|U}}\log\frac{\hat{Q}_{Z}}{\hat{Q}_{YZ}}$
$=\mathbb{E}_{P_{U}}\max_{\hat{Q}_{X}}\mathbb{E}_{P_{YZ|U}}\log\frac{1}{\hat{Q}_{Y|Z}}=\mathbb{E}_{P_{U}}\mathbb{E}_{P_{YZ|U}}\log\frac{1}{Q_{Y|Z}}<\infty.$
For the setting $\sigma_{Q,Z|X}\leq\sigma_{Q,Y|X}$, we would like to prove that there is a maximizer $P_{U|X}$ in which $U$ and $X$ are jointly Gaussian random variables.
Let
$\sigma^{2}_{X}$ $=\mathsf{Var}_{P}[X],$
$\sigma^{2}_{X|Y}$ $=\mathbb{E}[\mathsf{Var}_{P}[X|Y]],$
$\sigma^{2}_{X|Z}$ $=\mathbb{E}[\mathsf{Var}_{P}[X|Z]].$
Then, using the fact that $(X,Y,Z)$ is jointly Gaussian under $P$, we can write the constraints as
$h(X|YU)$ $\geq h(X|Y)-R=\frac{1}{2}\log(2\pi e\sigma_{X|Y}^{2})-R,$
$h(X|ZU)$ $\geq h(X|Z)-R=\frac{1}{2}\log(2\pi e\sigma_{X|Z}^{2})-R.$
Consider the following optimization problem where we relax the Gaussian assumption on the marginal distribution of $X$ under $P$:
$\sup_{P_{UX}}\mathbb{E}_{P_{U}}\left[\max_{\hat{Q}_{X}}D(P_{Y|U}\|\hat{Q}_{Y})-D(P_{Z|U}\|\hat{Q}_{Z})\right]$
where the supremum is over $P_{UX}$ satisfying
$h(X|YU)$ $\geq\frac{1}{2}\log(2\pi e\sigma_{X|Y}^{2})-R,$ (24)
$h(X|ZU)$ $\geq\frac{1}{2}\log(2\pi e\sigma_{X|Z}^{2})-R,$ (25)
and satisfying the following covariance constraint:
$\mathsf{Var}_{P}[X]\leq\sigma_{X}^{2}.$ (26)
In the above relaxed problem, the channels $P_{YZ|X}$ and $Q_{YZ|X}$ are kept as the same Gaussian channels. The relaxed problem is similar to channel coding type optimizations, as the distribution on $X$ is arbitrary. The advantage of the relaxation is that we can use Caratheodery’s theorem to restrict the alphabet size of random variable $U$ to $|\mathcal{U}|\leq 3$. This restriction is crucial when we employ the approach in [23] to prove the optimality of the Gaussian channel $P_{U|X}$. Herein, we also utilize the idea of "perturbated expressions" introduced in [21].
Let $G$ be a standard Gaussian random variable, independent of $(X,Y,Z)$, and (under either hypothesis) let
$\tilde{X}$ $=X+\delta\cdot G$
$\tilde{Y}$ $=Y+\delta\cdot G$
$\tilde{Z}$ $=Z+\delta\cdot G$
where $\delta>0$ is a small real number. Observe that the following Markov chain holds for any $P_{UX}$:
$U\rightarrow X\rightarrow\tilde{X}\rightarrow(\tilde{Y},\tilde{Z}).$
Consider the following perturbed optimization problem:
$\sup_{P_{UX}}\mathbb{E}_{P_{U}}\left[\max_{\hat{Q}_{X}}D(P_{\tilde{Y}|U}\|\hat{Q}_{Y})-D(P_{\tilde{Z}|U}\|\hat{Q}_{Z})\right]$
$-\delta\cdot h_{P}(\tilde{Y}|U)$ (27)
where the supremum is over $P_{UX}$ satisfying
$h_{P}(\tilde{X}|\tilde{Y}U)$ $\geq\frac{1}{2}\log(2\pi e\sigma_{X|Y}^{2})-R,$ (28)
$h_{P}(\tilde{X}|\tilde{Z}U)$ $\geq\frac{1}{2}\log(2\pi e\sigma_{X|Z}^{2})-R.$ (29)
and (26). Letting $\delta$ go to zero yields the original optimization problem. Therefore, it suffices to show the existence of a maximizer in which $U$ and $X$ are jointly Gaussian for the perturbed problem for any $\delta>0$.
Let $p^{*}(u,x,\tilde{x},\tilde{y},\tilde{z})$ be a maximizer of the optimization problem. [^3] Existence of a maximizer follows from the standard arguments in Appendix II of [23]. The Gaussian noise $G$ is considered to make sure that $\tilde{X}$ has a bounded density in order to get the proper convergence results. Alternatively, one can also utilize the result in [26] which does not necessitate showing the existence of a maximizer. Let
$(U^{*}_{1},X^{*}_{1},{\tilde{X}}^{*}_{1},\tilde{Y}^{*}_{1},\tilde{Z}^{*}_{1})$
and
$(U^{*}_{2},X^{*}_{2},{\tilde{X}}^{*}_{2},\tilde{Y}^{*}_{2},\tilde{Z}^{*}_{2})$
to be two independent copies of a maximizer $p^{*}(u,x,\tilde{x},\tilde{y},\tilde{z})$ of the optimization problem subject to the constraints in the problem. Further, let $V$ denote the optimal value. Let
$(U^{\prime},X^{\prime},\tilde{X}^{\prime},\tilde{Y}^{\prime},\tilde{Z}^{\prime})$
$:=((U^{*}_{1},U^{*}_{2}),(X^{*}_{1},X^{*}_{2}),(\tilde{X}^{*}_{1},\tilde{X}^{*}_{2}),(\tilde{Y}^{*}_{1},\tilde{Y}^{*}_{2}),(\tilde{Z}^{*}_{1},\tilde{Z}^{*}_{2}))$
and
$\begin{pmatrix}X_{+}\\ X_{-}\end{pmatrix}:=\begin{pmatrix}\sqrt{t}&\sqrt{1-t}\\ -\sqrt{1-t}&\sqrt{t}\end{pmatrix}\begin{pmatrix}X^{*}_{1}\\ X^{*}_{2}\end{pmatrix}$
where $t=\frac{1}{2}$. We define channel outputs $\tilde{X}_{+},\tilde{X}_{-},\allowbreak\tilde{Y}_{+},\tilde{Y}_{-},\tilde{Z}_{+},\tilde{Z}_{-}$ in the standard manner as in [23].
Set
$\hat{U}_{-}$ $:=(U^{\prime},\tilde{Y}_{+})$ $\quad\hat{U}_{+}$ $:=(U^{\prime},\tilde{Z}_{-}),$ (30)
Since rotations preserve mutual information, note that standard manipulations yield
$2V$
$=\mathbb{E}_{P_{U^{\prime}}}\Big{[}\max_{\hat{Q}_{X_{+}X_{-}}}D(P_{\tilde{Y}_{+}\tilde{Y}_{-}|U^{\prime}}\|\hat{Q}_{Y_{+}Y_{-}})$ (31)
$\quad-D(P_{\tilde{Z}_{+}\tilde{Z}_{-}|U^{\prime}}\|\hat{Q}_{Z_{+}Z_{-}})\Big{]}-\delta h_{P}(\tilde{Y}_{+}\tilde{Y}_{-}|U^{\prime})$
$=\mathbb{E}_{P_{U^{\prime}}}\bigg{[}\max_{\hat{Q}_{X_{+}X_{-}}}D(P_{\tilde{Y}_{+}\tilde{Y}_{-}|U^{\prime}}\|\hat{Q}_{\tilde{Y}_{+}\tilde{Y}_{-}})-D(P_{\tilde{Y}_{+}\tilde{Z}_{-}|U^{\prime}}\|\hat{Q}_{\tilde{Y}_{+}\tilde{Z}_{-}})$
$\quad+D(P_{\tilde{Y}_{+}\tilde{Z}_{-}|U^{\prime}}\|\hat{Q}_{\tilde{Y}_{+}\tilde{Z}_{-}})-D(P_{\tilde{Z}_{+}\tilde{Z}_{-}|U^{\prime}}\|\hat{Q}_{\tilde{Z}_{+}\tilde{Z}_{-}})\bigg{]}$
$\quad-\delta h_{P}(\tilde{Y}_{+}|U^{\prime})-\delta h_{P}(\tilde{Y}_{-}|\tilde{Y}_{+}U^{\prime})$
$\leq\mathbb{E}_{P_{U^{\prime}}}\bigg{[}\max_{\hat{Q}_{X_{+}X_{-}}}D(P_{\tilde{Y}_{-}|U^{\prime}\tilde{Y}_{+}}\|\hat{Q}_{\tilde{Y}_{-}|\tilde{Y}_{+}}|P_{\tilde{Y}_{+}|U^{\prime}})$ (32)
$\quad-D(P_{\tilde{Z}_{-}|U^{\prime}\tilde{Y}_{+}}\|\hat{Q}_{\tilde{Z}_{-}|\tilde{Y}_{+}}|P_{\tilde{Y}_{+}|U^{\prime}})\bigg{]}$
$\quad+\mathbb{E}_{P_{U^{\prime}}}\bigg{[}\max_{\hat{Q}_{X_{+}X_{-}}}D(P_{\tilde{Y}_{+}|U^{\prime}\tilde{Z}_{-}}\|\hat{Q}_{\tilde{Y}_{+}|\tilde{Z}_{-}}|P_{\tilde{Z}_{-}|U^{\prime}})$ (33)
$\quad-D(P_{\tilde{Z}_{+}|\tilde{Z}_{-}U^{\prime}}\|\hat{Q}_{\tilde{Z}_{+}|\tilde{Z}_{-}}|P_{\tilde{Z}_{-}|U^{\prime}})\bigg{]}$
$\quad-\delta h_{P}(\tilde{Y}_{+}|U^{\prime}\tilde{Z}_{-})-\delta h_{P}(\tilde{Y}_{-}|\tilde{Y}_{+}U^{\prime})-\delta I_{P}(\tilde{Y}_{+};\tilde{Z}_{-}|U^{\prime})$
$\leq\mathbb{E}_{P_{U^{\prime}\tilde{Y}_{+}}}\bigg{[}\max_{\hat{Q}_{X_{-}}}D(P_{\tilde{Y}_{-}|U^{\prime}\tilde{Y}_{+}}\|\hat{Q}_{\tilde{Y}_{-}})-D(P_{\tilde{Z}_{-}|U^{\prime}\tilde{Y}_{+}}\|\hat{Q}_{\tilde{Z}_{-}})\bigg{]}$
$\quad+\mathbb{E}_{P_{U^{\prime}\tilde{Z}_{-}}}\bigg{[}\max_{\hat{Q}_{X_{+}}}D(P_{\tilde{Y}_{+}|U^{\prime}\tilde{Z}_{-}}\|\hat{Q}_{\tilde{Y}_{+}})-D(P_{\tilde{Z}_{+}|\tilde{Z}_{-}U^{\prime}}\|\hat{Q}_{\tilde{Z}_{+}})\bigg{]}$
$\quad-\delta h_{P}(\tilde{Y}_{+}|U^{\prime}\tilde{Z}_{-})-\delta h_{P}(\tilde{Y}_{-}|\tilde{Y}_{+}U^{\prime})-\delta I_{P}(\tilde{Y}_{+};\tilde{Z}_{-}|U^{\prime})$
$=\mathbb{E}_{P_{\hat{U}^{\prime}_{-}}}\bigg{[}\max_{\hat{Q}_{X_{-}}}D(P_{\tilde{Y}_{-}|\hat{U}^{\prime}_{-}}\|\hat{Q}_{\tilde{Y}_{-}})-D(P_{\tilde{Z}_{-}|\hat{U}^{\prime}_{-}}\|\hat{Q}_{\tilde{Z}_{-}})\bigg{]}$
$+\mathbb{E}_{P_{\hat{U}^{\prime}_{+}}}\bigg{[}\max_{\hat{Q}_{X_{+}}}D(P_{\tilde{Y}_{+}|\hat{U}^{\prime}_{+}}\|\hat{Q}_{\tilde{Y}_{+}})-D(P_{\tilde{Z}_{+}|\hat{U}^{\prime}_{+}}\|\hat{Q}_{\tilde{Z}_{+}})\bigg{]}$
$\qquad-\delta h_{P}(\tilde{Y}_{+}|U^{\prime}_{+})-\delta h_{P}(\tilde{Y}_{-}|U^{\prime}_{-})-\delta I_{P}(\tilde{Y}_{+};\tilde{Z}_{-}|U^{\prime}).$ (34)
Moreover,
$h_{P}(\tilde{X}_{+}\tilde{X}_{-}|\tilde{Y}_{+}\tilde{Y}_{-}U^{\prime})$
$=h_{P}(\tilde{X}_{-}|\tilde{Y}_{-}U^{\prime}\tilde{Y}_{+})+h_{P}(\tilde{X}_{+}|\tilde{X}_{-}\tilde{Y}_{+}\tilde{Y}_{-}U^{\prime})$
$=h_{P}(\tilde{X}_{-}|\tilde{Y}_{-}U^{\prime}\tilde{Y}_{+})+h_{P}(\tilde{X}_{+}|\tilde{X}_{-}\tilde{Z}_{-}\tilde{Y}_{+}\tilde{Y}_{-}U^{\prime})$
$\leq h_{P}(\tilde{X}_{-}|\tilde{Y}_{-}U^{\prime}\tilde{Y}_{+})+h_{P}(\tilde{X}_{+}|\tilde{Y}_{+}U^{\prime}\tilde{Z}_{-})$
$=h_{P}(\tilde{X}_{-}|\tilde{Y}_{-}\hat{U}_{-})+h_{P}(\tilde{X}_{+}|\tilde{Y}_{+}\hat{U}_{+}).$
Similarly,
$h_{P}(\tilde{X}_{+}\tilde{X}_{-}|\tilde{Z}_{+}\tilde{Z}_{-}U^{\prime})$ $\leq h_{P}(\tilde{X}_{-}|\tilde{Z}_{-}\hat{U}_{-})+h_{P}(\tilde{X}_{+}|\tilde{Z}_{+}\hat{U}_{+}).$
[…]
$U^{\prime\prime}=(S,\hat{U}_{+}),~{}~{}~{}X^{\prime\prime}=X_{+},~{}~{}~{}\tilde{X}^{\prime\prime}=\tilde{X}_{+},$
$\tilde{Y}^{\prime\prime}=\tilde{Y}_{+},~{}~{}~{}\tilde{Z}^{\prime\prime}=\tilde{Z}_{+}.$
When $S=-$, we set
$U^{\prime\prime}=(S,\hat{U}_{-}),~{}~{}~{}X^{\prime\prime}=X_{-},~{}~{}~{}\tilde{X}^{\prime\prime}=\tilde{X}_{-},$
$\tilde{Y}^{\prime\prime}=\tilde{Y}_{-},~{}~{}~{}\tilde{Z}^{\prime\prime}=\tilde{Z}_{-}.$
Notice that the distribution on $U^{\prime\prime}X^{\prime\prime}$ is a candidate maximizer of the expression and hence must induce a value of at most $V$. Thus, from (34) we obtain that
$I_{P}(\tilde{Y}_{+};\tilde{Z}_{-}|U^{\prime})=0.$
The rest of the arguments mimic ones in [23] in Propositions 2, 8, and Corollary 3 of [23], and we omit the details. Using Proposition 2 in [23], we deduce that $I_{P}(\tilde{X}_{-};\tilde{X}_{+}|U^{\prime})=0$. By the Skitovic-Darmois characterization of Gaussians, we deduce that conditioned on $U^{\prime}$, random variables $X^{*}_{1},X^{*}_{2}$ are also jointly Gaussians, and they have the same covariance matrix (independent of $U^{\prime}$). Thus, a Gaussian $p(x|u)$ with the same covariance for each $u$ is an optimal solution of the optimization problem.